# MixLLM: LLM Quantization with Global Mixed-precision between Output-features and Highly-efficient System Design

### v1

@@ 4.3 System Performance
We have evaluated MixLLM for the single linear layer of token number ranging from 1 to 1024 with in features 4096 and out features 4096/14336, and compared it with the SOTA W4A16 (TRT-LLM) and QoQ [27], shown in Fig.5. It also shows MixLLM kernel performance of different percent of 8-bits (W4A8 0% 8-bit, W4.4A8 10% 8-bit, and W8A8 100% 8-bit). It shows that:
  • MixLLM outperforms the float16 counterpart for all token numbers, achieving 1.90$\times$, 2.75$\times$, and 1.88$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively.
  • MixLLM outperforms the SOTA W4A16 solution, achieving 1.26$\times$, 1.78$\times$, and 1.25$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively.
  • MixLLM achieves similar performance with QoQ with similar bit-width, achieving 0.99$\times$, 1.39$\times$, and 0.99$\times$ averaged speedup with MixLLM W4A8, W8A8, and W4.8A8 respectively. Note that MixLLM has better accuracy than QoQ (Tab.1, Tab.3).
Figure 5: The speedup of two types of single linear layers over torch float16 baseline on the A100 GPU.
4.4 Downstream Tasks Evaluation
@@
Tab.3 shows the accuracy of the downstream tasks on three popular LLMs. The result shows that:
  • MixLLM W4.4A8 outperforms all the 4-bit weight quantizations, with only 10% more weight consumption. For example, for the MMLU-Pro task, the average metric of MixLLM W4.4A8 is improved by 1.69, 6.93, and 0.93 over QoQ, QuaRot W4A4, and QuaRot W4A8, respectively.
  • MixLLM W8A8 is nearly lossless, showing higher accuracy than SmoothQuant. This comes from the group-wise quantized activation of MixLLM.
4.5 Detailed Analysis
4.5.1 Ablation Study
@@
4.5.2 High Precision Distribution
layer q_proj k_proj v_proj o_proj gate_proj up_proj down_proj
avg 8-bit percent (%) 3.93 12.36 71.22 18.70 0.73 1.46 53.82
4.5.3 One-pass vs. Progressive Search
As described in Sec.3.2, MixLLM searches the high-salience features within a single pass. We have tried the progressive procedure on Llama 3.1 8B and Mistral 7B models, which identifies smaller portions of the high-salience features iteratively. Results show that the accuracy is the same to the one-pass method to two decimal places. However, the progressive method shows much higher search time due to the repeated procedure. The one-pass method takes 7 minutes for each of the two models to search 10% high-salience features, while the progressive method that searches 2% high-salience iteratively takes 30 minutes to find top 10% features.
4.5.4 Overhead of Global Precision Search
Table 5: The overhead of global precision search in MixLLM.
@@ 4.5.4 Overhead of Global Precision Search
Tab.5 shows the global precision search overhead described in Sec.3.2. As noted in Sec.4.1, the calibration dataset has 128 samples with sequence length of 2048. We use a single A100 GPU for the 1.5B, 7B and 8B models, and 4 A100 GPUs for the 70B models to perform the search. We make use of device_map in huggingface for multi-GPU execution, which is sequential execution of layers on different devices. The 7B/8B models require about 7 minutes and the 70B models require less than 60 minutes to complete the search. Considering that the quantization only needs to be performed once, the searching algorithm is practical for the real workloads.
@@ 5 Summary
We have presented MixLLM, achieving high accuracy with low memory consumption and high system efficiency with the rarely explored optimization space of mixed-precision quantization between output features. MixLLM identifies the salience of each output feature according to the loss distance estimation w.r.t. the global model loss rather than local layer loss. By assigning larger bit-width to the features need it most, MixLLM achieves the superior accuracy to SOTA with low memory consumption. The sub-problems of different bit-widths are disjoint and can run in parallel efficiently on the GPU. We have identified the sweet spot of the quantizaiton configuration that is friendly to both accuracy and system efficiency. To address the challenge of system efficiency, we design the two-step dequantization to enable using int8 Tensor Core computation and the fast integer-float conversion to reduce the dequantization overhead. We have designed the end-to-end software pipeline to overlap the memory access, the dequantization computation with SIMT Core and the MatMul with Tensor Core. Experiment results show that MixLLM achieves superior accuracy to existing works and state-of-the-art system efficiency with low memory cost.

### v2

@@ 4.3 System Performance
4.3.1 Performance on A100 GPU
[…]
Figure 6: Speedup of MixLLM W8A8 of Qwen 2.5 7B over float16 baseline on A100 GPU.
[…]
4.3.2 Performance on H100 GPU
Figure 7: The speedup of two types of single linear layers over W4A16 baseline on H100 GPU.
We evaluate the performance of single linear layers on the NVIDIA H100 GPU, comparing the MixLLM W4A8 kernel against the W4A16 baseline (Fig. 7). To the best of our knowledge, an H100-optimized implementation of the QoQ solution Lin et al. (2025) is currently unavailable. The W4A16 baseline was established by identifying the optimal configuration through exhaustive tuning of the official CUTLASS group-wise and asymmetric-quantized W4A16 examples. Our analysis considers both a naive 4-bit layout and an advanced shuffled 4-bit layout. Results indicate that MixLLM consistently outperforms the W4A16 solutions across all configurations. Specifically, for an output feature size of 4096, MixLLM achieves average speedups of $1.81\times$ and $1.39\times$ over the naive and shuffled layouts, respectively; for a size of 14336, these speedups are $1.91\times$ and $1.34\times$.
@@ 4.3.2 Performance on H100 GPU
4.4 Downstream Tasks Evaluation
We first evaluate GSM8K Cobbe et al. (2021) on Qwen2.5 7B model, to validate the quantization efficacy on long-reasoning tasks. Using MixLLM W4.4A8 quantization, the strict-match metric only drops from 0.8 (the float16 model) to 0.792, only a 0.008 drop. Besides this, we evaluate a large number of the downstream tasks on three popular LLMs, shown in Tab.2. The result shows that: 1) MixLLM W4.4A8 outperforms all the 4-bit weight quantizations, with only 10% more bit-width. For example, for the MMLU-Pro task, the average metric of MixLLM W4.4A8 is improved by 1.69, 6.93, and 0.93 over QServe, QuaRot W4A4, and QuaRot W4A8, respectively. 2) MixLLM W8A8 is nearly lossless, showing higher accuracy than SmoothQuant. This comes from the group-wise quantized activation of MixLLM.
4.5 Ablation Study
@@
Model FP16
GPTQ
W4A16
AWQ
W4A16
SqueezeLLM
W4A16 0.45%
OmniQuant
W4A16/W4A4
AfineQuant
W4A16/W4A4
Atom
W4A4 128 outliers
SpinQuant
W4A8
MixLLM
W4.4A8
LLaMA 2 7B 5.47 5.59 5.60 5.57 5.58/14.26 5.58/12.69 6.03 5.7 5.55
LLaMA 3 8B 6.14 6.46 6.55 - - - 7.57 6.5 6.32
4.6 Comparison with More Related Work
Table 4: PPL of Atom and MixLLM with similar bit-width.
Models Llama 2 7B Llama 2 13B
Atom W4.4A8 5.64 5.03
MixLLM W4.4A8 5.54 4.93
Comparison with Atom. Tab.4 shows the perplexity of Atom and MixLLM with the similar bit-width (i.e., W4.4A8). We use 512 outliers in Atom so its weight is around 4.4-bit on average (using Atom’s opensourced code with git commit 7e3618b). It shows that MixLLM has much better accuracy than Atom. We also evaluated the kernel performance of Atom, showing that our W4.4A8 kernel achieves 1.56x speedup than Atom’s W4A8 kernel for the linear layer with sequence length 1024. This means our proposed memory and computation pipeline has much better performance than the related mixed-precision work.
Comparison with SliM-LLM. Tab.5 shows the perplexity of SliM-LLM W4A16 (it only supports weight-only) and MixLLM with different bit-width. SliM-LLM W4A16 is evaluated using the opensourced code. It is the mix of 3/4/5-bits in its code, and the precision is searched within each single layer, rather than using the global precision search. Note that MixLLM can support the mix of any bit-width. We evaluated MixLLM with the mix of 4-bit/6-bit (W4.2A8, 4.2-bit on average with 90% 4-bit and 10% 6-bit), and the mix of 3-bit/4-bit/6-bit (W3.9A8, 3.9-bit on average with 30% 3-bit, 60% 4-bit, and 10% 6-bit) in this table. We notice even MixLLM W4A8 (uniform 4-bit for the weight) can defeat SliM-LLM W4A16. This is understandable because SliM-LLM itself focuses on 2-bit and 3-bit optimization, and its paper’s main body does not show any 4-bit results. In contrast, MixLLM W3.9A8 can defeat MixLLM W4A8. This demonstrates that MixLLM’s global precision search can better allocate bit-width to the important weight elements to achieve the nearly lossless results. Note that the percentage of bit-widths is determined intuitively here, so that it may not be the optimal percentage to achieve the optimal accuracy. How to determine the portion of different bit-width can be a new research problem.
Table 5: PPL (wikitext2) of SliM-LLM and MixLLM.
Models Llama Qwen2.5
3.2 1B 3.1 8B 0.5B 1.5B 7B
SliM-LLM W4A16 10.80 6.55 14.85 9.68 7.02
MixLLM W3.9A8 10.20 6.50 13.68 9.51 6.97
MixLLM W4A8 (uniform) 10.36 6.54 14.43 9.66 7.03
MixLLM W4.2A8 10.07 6.43 13.53 9.45 6.93
MixLLM W4.4A8 10.05 6.42 13.42 9.44 6.92
We compare MixLLM with more recent quantization works according to the reported numbers in their papers (Tab.3), showing that MixLLM achieves superior accuracy to a broad range of related works with similar memory consumption.
4.7 Overhead of Global Precision Search
Table 6: The overhead of global precision search.
@@ 4.7 Overhead of Global Precision Search
Tab.6 shows the global precision search overhead described in Sec.3.2. As noted in Sec.4.1, the calibration dataset has 128 samples with sequence length of 2048. We use a single A100 GPU for the 1.5B, 7B and 8B models, and 4 A100 GPUs for the 70B models to perform the search. We make use of device_map in huggingface for multi-GPU execution, which is sequential execution of layers on different devices. The 7B/8B models require about 7 minutes and the 70B models require less than 60 minutes to complete the search. Considering that the quantization only needs to be performed once, the searching algorithm is practical for the real workloads. In contrast, SliM-LLM takes more than 3 hours to determine the precision of Llama 3.1 8B model and more than 4 hours for Qwen2.5 7B model according to our experiments.
4.8 High Precision Distribution
layer (xx_proj) q k v o gate up down
avg 8-bit (%) 3.93 12.36 71.22 18.70 0.73 1.46 53.82
4.9 One-pass vs. Progressive Search
As described in Sec.3.2, MixLLM searches the high-salience features within a single pass rather than iteratively identifying the high-salience parts in a smaller step, as we observe the single-pass method show similar results with the iterative method and saves a lot of computation overhead than the latter. We have tried the progressive procedure on Llama 3.1 8B and Mistral 7B models, which identifies smaller portions of the high-salience features iteratively. Results show that the accuracy is the same to the one-pass method to two decimal places. However, the progressive method shows much higher search time due to the repeated procedure. The one-pass method takes 7 minutes for each of the two models to search 10% high-salience features, while the progressive method that searches 2% high-salience iteratively takes 30 minutes to find top 10% features.
4.10 Working with KV Quantization
Table 8: PPL of enabling KV Quantization in MixLLM.
Models Llama 3.1 8B Qwen2.5 7B
MixLLM W4.4A8 w/o KV quant 6.42 6.92
MixLLM W4.4A8 KV8 6.42 6.96
It is straightforward to apply any KV quantization technology together with MixLLM’s weight-activation quantization, as KV quantization is orthogonal to the weight-activation quantization. Tab.8 shows the perplexity (wikitext2) of Llama 3.1 8B and Qwen 2.5 7B for enabling and disabling KV quantization. It shows that the 8-bit KV quantization is nearly lossless upon MixLLM.
4.11 Calibration Dataset for Salience Search
We use 128 samples for the calibration to search the high-salience channels in MixLLM, which is a common configuration in the existing solutions. We also find that a smaller dataset can still identify the high-salience channels in MixLLM. The perplexity of using different samples for MixLLM salience search are shown in Tab.9. This shows that shrinking the sample size from 128 to 16 for the global salience search has negligible effect to the accuracy.
Table 9: PPL (wikitext2) when using different number of samples for salience search (W4.4A8 quantized).
#samples 128 64 32 16
Llama 3.1 8B 6.42 6.42 6.42 6.42
Qwen2.5 7B 6.92 6.92 6.92 6.93
We also use different dataset families between salience searching and perplexity evaluation to evaluate the effect of input distribution shift. When using c4 dataset for salience search and evaluate the perplexity on wikitext2 (W4.4A8 quantized), the Llama 3.1 8B and Qwen2.5 7B has the same perplexity with that using wikitext2 for the calibration. This shows that changing the calibration dataset does not affect the accuracy for the same task. In another word, when the data distribution between the calibration dataset and real input is different, MixLLM’s salience search algorithm still works well.
@@ 5 Summary
We have presented MixLLM, achieving high accuracy with low memory consumption and high system efficiency with the rarely explored optimization space of mixed-precision quantization between output features. MixLLM identifies the salience of each output feature according to the loss w.r.t. the global model rather than each single layer. By assigning larger bit-width to the features need it the most, MixLLM achieves the superior accuracy to SOTA with low memory consumption. The sub-problems of different bit-widths are disjoint and can run in parallel efficiently on the GPU. We have identified the sweet spot of the quantization configuration that is friendly to both accuracy and system efficiency. To address the challenge of system efficiency, we design the two-step dequantization to enable using int8 Tensor Core and the fast integer-float conversion to reduce the dequantization overhead. We have designed the end-to-end software pipeline to overlap the memory access, the dequantization computation with SIMT Core and the MatMul with Tensor Core to the best. Experiment results show that MixLLM achieves superior accuracy to existing works and state-of-the-art system efficiency with low memory cost.